# Statistical-Computational Trade-offs for Recursive Adaptive Partitioning Estimators

### v2

@@ Appendix B Length of Query Paths
First, we use Theorem 13.5 in Wainwright (2019) with the choice $t=\max\left\{\epsilon_{n},\frac{2\sigma^{2}\log n}{n\epsilon_{n}}\right\}$ to obtain a probability at least $1-1/n$ event on which
$\begin{split}\left\|\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n})-f^{*}\right\|_{n}^{2}&\leq 16t\epsilon_{n}\\ &=\max\left\{16\epsilon_{n}^{2},\frac{32\sigma^{2}\log n}{n}\right\}\\ &\lesssim\frac{\sigma^{2}}{n}\max\left\{2^{s}\log d,\log n\right\}.\end{split}$ (E.26)
Likewise, we use Theorem 14.12 in Wainwright (2019) with the choice $\delta=\max\left\{\delta_{n},\sqrt{\frac{4M^{2}\log n}{c_{1}n}}\right\}$ to obtain a probability at least $1-1/n$ event on which
$\begin{split}R\left(\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n}),f^{*}\right)&\leq 2\left\|\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n})-f^{*}\right\|_{n}^{2}+\delta^{2}\\ &\lesssim\frac{\sigma^{2}+M^{2}}{n}\max\left\{2^{s}\log d,\log n\right\}.\end{split}$ (E.27)
Let $\mathcal{E}$ be the event on which (E.27) holds. We then write
$\mathfrak{R}\left(\hat{f}_{\operatorname{ERM}},f^{*}_{0},d,n\right)=\mathbb{E}\left\{R\left(\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n}),f^{*}\right)\mathbb{1}_{\mathcal{E}}\right\}+\mathbb{E}\left\{R\left(\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n}),f^{*}\right)\mathbb{1}_{\mathcal{E}^{c}}\right\}.$ (E.28)
@@
$\|\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n})\|_{\infty}\leq\max_{1\leq i\leq n}\left|Y_{i}\right|,$ (E.29)
@@
$R\left(\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n}),f^{*}\right)\leq\left(\max_{1\leq i\leq n}\left|Y_{i}\right|+\left\|f^{*}\right\|_{\infty}\right)^{2}.$ (E.30)
@@
$\max_{1\leq i\leq n}\left|Y_{i}\right|\lesssim\sigma\left(\sqrt{\log n}+t\right)$ (E.31)
with probability at least $1-2e^{-t^{2}}$, for any $t>0$. Plugging this into (E.30) and applying Cauchy-Schwarz twice gives
$R\left(\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n}),f^{*}\right)\lesssim\left\|f^{*}\right\|_{\infty}^{2}+\sigma^{2}\log n+\sigma^{2}u$ (E.32)
@@
$\mathbb{P}\left\{R\left(\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n}),f^{*}\right)\geq u\right\}\leq 2e^{-C(u-\|f^{*}\|_{\infty}^{2}+\sigma^{2}\log n)/\sigma^{2}}$ (E.33)
@@
[…]
@@
$\mathfrak{R}\left(\hat{f}_{\operatorname{ERM}},f^{*}_{0},d,n\right)\lesssim\frac{(\sigma^{2}+M^{2})(2^{s}\log d+\log n)}{n}$ (E.35)
@@
Appendix F Stable MSP Properties
Proposition F.1.
@@ Appendix F Stable MSP Properties
$\mathcal{C}=\left\{\mathbf{x}\in\{\pm 1\}^{d}\colon x_{j}=z_{j}~{}\forall j\in J(\mathcal{C})\right\}.$ (F.1)
@@
$\alpha_{S}^{\mathcal{C}}\coloneqq\sum_{U\in\mathcal{S},U\backslash J(\mathcal{C})=S}\left(\prod_{j\in U\cap J(\mathcal{C})}z_{j}\right)\alpha_{U},$ (F.2)
@@
$\mathcal{S}|_{\mathcal{C}}\coloneqq\left\{S\colon\alpha_{S}^{\mathcal{C}}\neq 0\right\}.$ (F.3)
Define the map $\Phi\colon S\mapsto S\backslash J(\mathcal{C})$ and observe that $\mathcal{S}|_{\mathcal{C}}\subset\Phi(\mathcal{S})$. We first claim that $\Phi(\mathcal{S})$ is MSP. To see this, simply note that for any $i$, we have
$\Phi(S_{i})\backslash\cup_{j=1}^{i-1}\Phi(S_{j})=\left(S_{i}\backslash\cup_{j=1}^{i-1}S_{j}\right)\backslash J(\mathcal{C}),$ (F.4)
@@
$\bigcup_{T\in\Phi(\mathcal{S})}\left\{\sum_{U\in\Phi^{-1}(T)}\left(\prod_{j\in U\cap J(\mathcal{C})}z_{j}\right)\alpha_{U}=0\right\}.$ (F.5)
@@
We shall prove the statement by induction on the number of covariates $d$, with the base case $d=0$ trivial. Now suppose the equivalence is true up to some value $d-1$, and consider $f^{*}\colon\{\pm 1\}^{d}\to\mathbb{R}$. Assume first that for any $\mathcal{C}\subset\{\pm 1\}^{d}$, $f^{*}|_{\mathcal{C}}$ is MSP. Then for any $\mathcal{C}\neq\{\pm 1\}^{d}$, by the inductive hypothesis, we have that
$\max_{j\in S^{*}\backslash J(\mathcal{C})}\operatorname{Corr}^{2}\left\{X_{j},f^{*}(\mathbf{X})~{}|~{}\mathbf{X}\in\mathcal{C}\right\}>0.$ (F.6)
@@
$\operatorname{Corr}\left\{X_{j},f^{*}(\mathbf{X})~{}|~{}\mathbf{X}\in\mathcal{C}\right\}=\alpha_{j}^{\mathcal{C}}.$ (F.7)
@@
$\operatorname{Corr}^{2}\left\{X_{k},f^{*}(\mathbf{X})\right\}=\alpha_{k}^{2}>0,$ (F.8)
which completes the proof that $f^{*}$ satisfies SID. Conversely, suppose that $f^{*}$ satisfies SID. Then by the inductive hypothesis, we see that $f^{*}|_{\mathcal{C}}$ is MSP for any $\mathcal{C}\neq\{\pm 1\}^{d}$ and it remains to prove that $f^{*}$ itself is MSP. To see this, we first see that SID implies (F.8) for some $k\in[d]$, which in turn means that $\{k\}\in\mathcal{S}$. Set $\mathcal{C}=\left\{\mathbf{x}\in\{\pm 1\}^{d}\colon x_{k}=1\right\}$. Then since $f^{*}|_{\mathcal{C}}$ is MSP by the inductive hypothesis, there is an ordering of indices such
$\left|S_{i}^{\mathcal{C}}\backslash\cup_{j=1}^{i-1}S_{j}^{\mathcal{C}}\right|\leq 1$ (F.9)
for any $i=1,2,\ldots,r\coloneqq\left|\mathcal{S}|_{\mathcal{C}}\right|$. But then by the definition of $S^{\mathcal{C}}_{i}$, there exists some $S_{i}\in\mathcal{S}$ such that $S_{i}\backslash\left\{k\right\}=S^{\mathcal{C}}_{i}$, or in other words, that either $S_{i}^{\mathcal{C}}\in\mathcal{S}$ or $S_{i}^{\mathcal{C}}\cup\left\{k\right\}\in\mathcal{S}$. Augmenting the sequence with $S_{0}\coloneqq\{k\}$ gives
$\left|S_{i}\backslash\cup_{j=1}^{i-1}S_{j}\right|\leq 1$ (F.10)
@@
Appendix G Proof of Theorem 7.5
@@
$p_{k}(\mathcal{C})\coloneqq\frac{N(\mathcal{T}_{k,1}(\mathcal{C}))}{N(\mathcal{C})}.$ (G.1)
@@
$\hat{\Delta}(k;\mathcal{C},\mathcal{D}_{n})=p_{k}(\mathcal{C})(1-p_{k}(\mathcal{C}))\left(\bar{Y}_{\mathcal{T}_{k,1}(\mathcal{C})}-\bar{Y}_{\mathcal{T}_{k,-1}(\mathcal{C})}\right)^{2}.$ (G.2)
@@
$\Delta(k;\mathcal{C},\mathcal{D}_{n})\coloneqq\frac{1}{4}\left(\mathbb{E}\left\{Y~{}|~{}\mathbf{X}\in\mathcal{T}_{k,1}(\mathcal{C})\right\}-\mathbb{E}\left\{Y~{}|~{}\mathbf{X}\in\mathcal{T}_{k,-1}(\mathcal{C})\right\}\right)^{2}.$ (G.3)
@@
$\hat{\Delta}^{1/2}(k;\mathcal{C},\mathcal{D}_{n})\coloneqq\sqrt{p_{k}(\mathcal{C})(1-p_{k}(\mathcal{C}))}\left(\bar{Y}_{\mathcal{T}_{k,1}(\mathcal{C})}-\bar{Y}_{\mathcal{T}_{k,-1}(\mathcal{C})}\right).$ (G.4)
$\Delta^{1/2}(k;\mathcal{C},\mathcal{D}_{n})\coloneqq\frac{1}{2}\left(\mathbb{E}\left\{Y~{}|~{}\mathbf{X}\in\mathcal{T}_{k,1}(\mathcal{C})\right\}-\mathbb{E}\left\{Y~{}|~{}\mathbf{X}\in\mathcal{T}_{k,-1}(\mathcal{C})\right\}\right).$ (G.5)
@@
$\mathfrak{C}\coloneqq\left\{\mathcal{C}\subset\{\pm 1\}^{d}\colon J(\mathcal{C})\subset S^{*}\right\}.$ (G.6)
${\mathfrak{C}^{*}}\coloneqq\left\{\mathcal{C}\in\mathfrak{C}\colon f^{*}|_{\mathcal{C}}~{}\text{is not constant}\right\}.$ (G.7)
@@
$\mathbb{P}\left\{\sqrt{N(\mathcal{C})}\left|\hat{\Delta}^{1/2}(k;\mathcal{C},\mathcal{D}_{n})-\Delta^{1/2}(k;\mathcal{C},\mathcal{D}_{n})\right|\geq\epsilon\right\}\leq 6\exp\left(-\frac{\epsilon^{2}}{18(9\|f^{*}\|_{\infty}^{2}+\sigma^{2})}\right),$ (G.8)
$\mathbb{P}\left\{\sqrt{N(\mathcal{C})}\left|\bar{Y}_{\mathcal{C}}-\mathbb{E}\left\{Y~{}|~{}\mathbf{X}\in\mathcal{C}\right\}\right|\geq\epsilon\right\}\leq 2\exp\left(-\frac{\epsilon^{2}}{2(\|f^{*}\|_{\infty}^{2}+\sigma^{2})}\right),$ (G.9)
@@
$\mathbb{P}\left\{N(\mathcal{C})\leq\frac{(1-\epsilon)n}{2^{|J(\mathcal{C})|}}\right\}\leq\exp\left(-\frac{n\epsilon^{2}}{2^{|J(\mathcal{C})|+1}}\right).$ (G.10)
@@
First, notice that (G.10) is just an application of Chernoff’s inequality (Vershynin, 2018). Next, conditioned on $\mathbf{X}\in\mathcal{C}$, we see that $Y=f^{*}(\mathbf{X})+\epsilon$ has sub-Gaussian parameter $\sqrt{\|f^{*}\|_{\infty}^{2}+\sigma^{2}}$ (Wainwright, 2019). Equation G.9 thus follows immediately from Hoeffding’s inequality conditioning on $N(\mathcal{C})$.
Finally, to prove (G.8), we again condition on $N(\mathcal{C})$ and write
@@
$\mathbb{P}\left\{I+II\geq\frac{\epsilon}{3}\right\}\leq 4\exp\left(-\frac{N(\mathcal{C})\epsilon^{2}}{18\left(\left\|f^{*}\right\|_{\infty}^{2}+\sigma^{2}\right)}\right).$ (G.11)
@@
$\mathbb{P}\left\{III\geq\frac{\epsilon}{3}\right\}\leq 2\exp\left(-\frac{N(\mathcal{C})\epsilon^{2}}{162\left\|f^{*}\right\|_{\infty}^{2}}\right).$ (G.12)
Combining (G.11) and (G.12), dividing $\epsilon$ by $\sqrt{N(\mathcal{C})}$ and then deconditioning on $N(\mathcal{C})$, we complete the proof of (G.8). ∎
Lemma G.2.
For any $10<\delta<$ 1, there is an event with probability at least $1-\delta$ over which the following inequalities hold:
$\min_{\mathcal{C}\in{\mathfrak{C}^{*}}}\left(\max_{k\in S^{*}}\hat{\Delta}^{1/2}(k;\mathcal{C},\mathcal{D}_{n})-\sqrt{2^{|J(\mathcal{C})|}\lambda}+\sqrt{\frac{18(9\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+2)\log 3+\log(2/\delta)\right)}{N(\mathcal{C})}}\right)\geq 0,$ (G.13)
$\max_{\mathcal{C}\in\mathfrak{C}}\left(\max_{k\notin S^{*}}\hat{\Delta}^{1/2}(k;\mathcal{C},\mathcal{D}_{n})-\sqrt{\frac{18(9\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+2)\log 3+\log(2d/\delta)\right)}{N(\mathcal{C})}}\right)\leq 0,$ (G.14)
@@
$\max_{\mathcal{C}\in\mathfrak{C}}\left(\left|\bar{Y}_{I(\mathcal{C})}-\mathbb{E}\left\{Y~{}|~{}\mathbf{x}\in\mathcal{C}\right\}\right|-\sqrt{\frac{2(\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+1)\log 3+\log(2/\delta)\right)}{N(\mathcal{C})}}\right)\leq 0.$ (G.15)
@@
$\min_{\mathcal{C}\in\mathfrak{C}}\frac{N(\mathcal{C})2^{|J(\mathcal{C})|+1}}{n}\geq 1.$ (G.16)
@@
$\mathfrak{C}=\bigcup_{S\subset S^{*}}\left\{\mathcal{C}\colon J(\mathcal{C})=S\right\},$ (G.17)
@@
$\left|\mathfrak{C}\right|=\sum_{k=0}^{s}\binom{s}{k}2^{k}=3^{s}.$ (G.18)
Using (G.10) from Lemma G.1 together with a union bound, we have $N(\mathcal{C})\geq\frac{n}{2^{|J(\mathcal{C})|+1}}$ uniformly for all $\mathcal{C}\in{\mathfrak{C}^{*}}$ with probability at least $1-\exp\left(s\log 3-\frac{n}{2^{|J(\mathcal{C})|+3}}\right)$, which proves (G.16).
@@
$\operatorname{Corr}\left\{x_{k},f(\mathbf{x})~{}|~{}\mathbf{x}\in\mathcal{C}\right\}=\frac{1}{2}\left(\mathbb{E}\left\{Y~{}|~{}\mathbf{x}\in\mathcal{T}_{k,1}(\mathcal{C})\right\}-\mathbb{E}\left\{Y~{}|~{}\mathbf{x}\in\mathcal{T}_{k,-1}(\mathcal{C})\right\}\right)=\Delta^{1/2}(k;\mathcal{C},\mathcal{D}_{n}).$ (G.19)

### v3

@@ Appendix C Length of Query Paths
First, we use Theorem 13.5 in Wainwright (2019) with the choice $t=\max\left\{\epsilon_{n},\frac{2\sigma^{2}\log n}{n\epsilon_{n}}\right\}$ to obtain a probability at least 11/n1-1/n event on which
$\begin{split}\left\|\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n})-f^{*}\right\|_{n}^{2}&\leq 16t\epsilon_{n}\\ &=\max\left\{16\epsilon_{n}^{2},\frac{32\sigma^{2}\log n}{n}\right\}\\ &\lesssim\frac{\sigma^{2}}{n}\max\left\{2^{s}\log d,\log n\right\}.\end{split}$ (F.26)
Likewise, we use Theorem 14.12 in Wainwright (2019) with the choice $\delta=\max\left\{\delta_{n},\sqrt{\frac{4M^{2}\log n}{c_{1}n}}\right\}$ to obtain a probability at least 11/n1-1/n event on which
$\begin{split}R\left(\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n}),f^{*}\right)&\leq 2\left\|\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n})-f^{*}\right\|_{n}^{2}+\delta^{2}\\ &\lesssim\frac{\sigma^{2}+M^{2}}{n}\max\left\{2^{s}\log d,\log n\right\}.\end{split}$ (F.27)
Let $\mathcal{E}$ be the event on which (F.27) holds. We then write
$\mathfrak{R}\left(\hat{f}_{\operatorname{ERM}},f^{*}_{0},d,n\right)=\mathbb{E}\left\{R\left(\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n}),f^{*}\right)\mathbb{1}_{\mathcal{E}}\right\}+\mathbb{E}\left\{R\left(\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n}),f^{*}\right)\mathbb{1}_{\mathcal{E}^{c}}\right\}.$ (F.28)
@@
$\|\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n})\|_{\infty}\leq\max_{1\leq i\leq n}\left|Y_{i}\right|,$ (F.29)
@@
$R\left(\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n}),f^{*}\right)\leq\left(\max_{1\leq i\leq n}\left|Y_{i}\right|+\left\|f^{*}\right\|_{\infty}\right)^{2}.$ (F.30)
@@
$\max_{1\leq i\leq n}\left|Y_{i}\right|\lesssim\sigma\left(\sqrt{\log n}+t\right)$ (F.31)
with probability at least $1-2e^{-t^{2}}$, for any $t>0$. Plugging this into (F.30) and applying Cauchy-Schwarz twice gives
$R\left(\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n}),f^{*}\right)\lesssim\left\|f^{*}\right\|_{\infty}^{2}+\sigma^{2}\log n+\sigma^{2}u$ (F.32)
@@
$\mathbb{P}\left\{R\left(\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n}),f^{*}\right)\geq u\right\}\leq 2e^{-C(u-\|f^{*}\|_{\infty}^{2}+\sigma^{2}\log n)/\sigma^{2}}$ (F.33)
@@
$\begin{split}\mathbb{E}\left\{R\left(\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n}),f^{*}\right)\mathbb{1}_{\mathcal{E}^{c}}\right\}&=\int_{0}^{\infty}\mathbb{P}\left\{R\left(\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n}),f^{*}\right)\mathbb{1}_{\mathcal{E}^{c}}\geq u\right\}du\\ &\leq\int_{0}^{\infty}\min\left\{\mathbb{P}\left\{R\left(\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n}),f^{*}\right)\geq u\right\},\mathbb{P}\left\{\mathcal{E}^{c}\right\}\right\}du\\ &\leq\int_{0}^{\infty}\min\left\{\frac{2}{n},2e^{-C(u-\|f^{*}\|_{\infty}^{2}+\sigma^{2}\log n)/\sigma^{2}}\right\}du\\ &\lesssim\frac{\sigma^{2}\log n+\left\|f^{*}\right\|_{\infty}^{2}}{n}.\end{split}$ (F.34)
@@
$\mathfrak{R}\left(\hat{f}_{\operatorname{ERM}},f^{*}_{0},d,n\right)\lesssim\frac{(\sigma^{2}+M^{2})(2^{s}\log d+\log n)}{n}$ (F.35)
@@
Appendix G Stable MSP Properties
Proposition G.1.
@@ Appendix G Stable MSP Properties
$\mathcal{C}=\left\{\mathbf{x}\in\{\pm 1\}^{d}\colon x_{j}=z_{j}~\forall j\in J(\mathcal{C})\right\}.$ (G.1)
@@
[…]
@@
$\mathcal{S}|_{\mathcal{C}}\coloneqq\left\{S\colon\alpha_{S}^{\mathcal{C}}\neq 0\right\}.$ (G.3)
[…]
@@
$\bigcup_{T\in\Phi(\mathcal{S})}\left\{\sum_{U\in\Phi^{-1}(T)}\left(\prod_{j\in U\cap J(\mathcal{C})}z_{j}\right)\alpha_{U}=0\right\}.$ (G.5)
@@
We shall prove the statement by induction on the number of covariates $d$, with the base case $d=0$ trivial. Now suppose the equivalence is true up to some value d1d-1, and consider $f^{*}\colon\{\pm 1\}^{d}\to\mathbb{R}$. Assume first that for any $\mathcal{C}\subset\{\pm 1\}^{d}$, $f^{*}|_{\mathcal{C}}$ is MSP. Then for any $\mathcal{C}\neq\{\pm 1\}^{d}$, by the inductive hypothesis, we have that
$max_{j\in S^{*}\backslash J(\mathcal{C})}\operatorname{Corr}^{2}\left\{X_{j},f^{*}(\mathbf{X})~|~\mathbf{X}\in\mathcal{C}\right\}>0.$ (G.6)
@@
$\operatorname{Corr}\left\{X_{j},f^{*}(\mathbf{X})~|~\mathbf{X}\in\mathcal{C}\right\}=\alpha_{j}^{\mathcal{C}}.$ (G.7)
@@
$\operatorname{Corr}^{2}\left\{X_{k},f^{*}(\mathbf{X})\right\}=\alpha_{k}^{2}>0,$ (G.8)
which completes the proof that $f^{*}$ satisfies SID. Conversely, suppose that $f^{*}$ satisfies SID. Then by the inductive hypothesis, we see that $f^{*}|_{\mathcal{C}}$ is MSP for any $\mathcal{C}\neq\{\pm 1\}^{d}$ and it remains to prove that $f^{*}$ itself is MSP. To see this, we first see that SID implies (G.8) for some $k\in[d]$, which in turn means that $\{k\}\in\mathcal{S}$. Set $\mathcal{C}=\left\{\mathbf{x}\in\{\pm 1\}^{d}\colon x_{k}=1\right\}$. Then since $f^{*}|_{\mathcal{C}}$ is MSP by the inductive hypothesis, there is an ordering of indices such
[…]
@@
Appendix H Proof of Theorem 7.5
@@
$p_{k}(\mathcal{C})\coloneqq\frac{N(\mathcal{T}_{k,1}(\mathcal{C}))}{N(\mathcal{C})}.$ (H.1)
@@
$\hat{\Delta}(k;\mathcal{C},\mathcal{D}_{n})=p_{k}(\mathcal{C})(1-p_{k}(\mathcal{C}))\left(\bar{Y}_{\mathcal{T}_{k,1}(\mathcal{C})}-\bar{Y}_{\mathcal{T}_{k,-1}(\mathcal{C})}\right)^{2}.$ (H.2)
@@
$\Delta(k;\mathcal{C},\mathcal{D}_{n})\coloneqq\frac{1}{4}\left(\mathbb{E}\left\{Y~|~\mathbf{X}\in\mathcal{T}_{k,1}(\mathcal{C})\right\}-\mathbb{E}\left\{Y~|~\mathbf{X}\in\mathcal{T}_{k,-1}(\mathcal{C})\right\}\right)^{2}.$ (H.3)
@@
$\hat{\Delta}^{1/2}(k;\mathcal{C},\mathcal{D}_{n})\coloneqq\sqrt{p_{k}(\mathcal{C})(1-p_{k}(\mathcal{C}))}\left(\bar{Y}_{\mathcal{T}_{k,1}(\mathcal{C})}-\bar{Y}_{\mathcal{T}_{k,-1}(\mathcal{C})}\right).$ (H.4)
$\Delta^{1/2}(k;\mathcal{C},\mathcal{D}_{n})\coloneqq\frac{1}{2}\left(\mathbb{E}\left\{Y~|~\mathbf{X}\in\mathcal{T}_{k,1}(\mathcal{C})\right\}-\mathbb{E}\left\{Y~|~\mathbf{X}\in\mathcal{T}_{k,-1}(\mathcal{C})\right\}\right).$ (H.5)
@@
$\mathfrak{C}\coloneqq\left\{\mathcal{C}\subset\{\pm 1\}^{d}\colon J(\mathcal{C})\subset S^{*}\right\}.$ (H.6)
${\mathfrak{C}^{*}}\coloneqq\left\{\mathcal{C}\in\mathfrak{C}\colon f^{*}|_{\mathcal{C}}~\text{is not constant}\right\}.$ (H.7)
@@
$\mathbb{P}\left\{\sqrt{N(\mathcal{C})}\left|\hat{\Delta}^{1/2}(k;\mathcal{C},\mathcal{D}_{n})-\Delta^{1/2}(k;\mathcal{C},\mathcal{D}_{n})\right|\geq\epsilon\right\}\leq 6\exp\left(-\frac{\epsilon^{2}}{18(9\|f^{*}\|_{\infty}^{2}+\sigma^{2})}\right),$ (H.8)
$\mathbb{P}\left\{\sqrt{N(\mathcal{C})}\left|\bar{Y}_{\mathcal{C}}-\mathbb{E}\left\{Y~|~\mathbf{X}\in\mathcal{C}\right\}\right|\geq\epsilon\right\}\leq 2\exp\left(-\frac{\epsilon^{2}}{2(\|f^{*}\|_{\infty}^{2}+\sigma^{2})}\right),$ (H.9)
@@
$\mathbb{P}\left\{N(\mathcal{C})\leq\frac{(1-\epsilon)n}{2^{|J(\mathcal{C})|}}\right\}\leq\exp\left(-\frac{n\epsilon^{2}}{2^{|J(\mathcal{C})|+1}}\right).$ (H.10)
@@
First, notice that (H.10) is just an application of Chernoff’s inequality (Vershynin, 2018). Next, conditioned on $\mathbf{X}\in\mathcal{C}$, we see that $Y=f^{*}(\mathbf{X})+\epsilon$ has sub-Gaussian parameter $\sqrt{\|f^{*}\|_{\infty}^{2}+\sigma^{2}}$ (Wainwright, 2019). Equation H.9 thus follows immediately from Hoeffding’s inequality conditioning on $N(\mathcal{C})$.
Finally, to prove (H.8), we again condition on $N(\mathcal{C})$ and write
@@
$\mathbb{P}\left\{I+II\geq\frac{\epsilon}{3}\right\}\leq 4\exp\left(-\frac{N(\mathcal{C})\epsilon^{2}}{18\left(\left\|f^{*}\right\|_{\infty}^{2}+\sigma^{2}\right)}\right).$ (H.11)
@@
$\mathbb{P}\left\{III\geq\frac{\epsilon}{3}\right\}\leq 2\exp\left(-\frac{N(\mathcal{C})\epsilon^{2}}{162\left\|f^{*}\right\|_{\infty}^{2}}\right).$ (H.12)
Combining (H.11) and (H.12), dividing $\epsilon$ by $\sqrt{N(\mathcal{C})}$ and then deconditioning on $N(\mathcal{C})$, we complete the proof of (H.8). ∎
Lemma H.2.
For any $0<\delta<1$, there is an event with probability at least $1-\delta$ over which the following inequalities hold:
$\min_{\mathcal{C}\in{\mathfrak{C}^{*}}}\left(\max_{k\in S^{*}}\hat{\Delta}^{1/2}(k;\mathcal{C},\mathcal{D}_{n})-\sqrt{2^{|J(\mathcal{C})|}\lambda}+\sqrt{\frac{18(9\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+2)\log 3+\log(2/\delta)\right)}{N(\mathcal{C})}}\right)\geq 0,$ (H.13)
$\max_{\mathcal{C}\in\mathfrak{C}}\left(\max_{k\notin S^{*}}\hat{\Delta}^{1/2}(k;\mathcal{C},\mathcal{D}_{n})-\sqrt{\frac{18(9\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+2)\log 3+\log(2d/\delta)\right)}{N(\mathcal{C})}}\right)\leq 0,$ (H.14)
@@
$\max_{\mathcal{C}\in\mathfrak{C}}\left(\left|\bar{Y}_{I(\mathcal{C})}-\mathbb{E}\left\{Y~|~\mathbf{x}\in\mathcal{C}\right\}\right|-\sqrt{\frac{2(\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+1)\log 3+\log(2/\delta)\right)}{N(\mathcal{C})}}\right)\leq 0.$ (H.15)
@@
$\min_{\mathcal{C}\in\mathfrak{C}}\frac{N(\mathcal{C})2^{|J(\mathcal{C})|+1}}{n}\geq 1.$ (H.16)
@@
$\mathfrak{C}=\bigcup_{S\subset S^{*}}\left\{\mathcal{C}\colon J(\mathcal{C})=S\right\},$ (H.17)
@@
$\left|\mathfrak{C}\right|=\sum_{k=0}^{s}\binom{s}{k}2^{k}=3^{s}.$ (H.18)
Using (H.10) from Lemma H.1 together with a union bound, we have $N(\mathcal{C})\geq\frac{n}{2^{|J(\mathcal{C})|+1}}$ uniformly for all $\mathcal{C}\in{\mathfrak{C}^{*}}$ with probability at least $1-\exp\left(s\log 3-\frac{n}{2^{|J(\mathcal{C})|+3}}\right)$, which proves (H.16).
@@
$\operatorname{Corr}\left\{x_{k},f(\mathbf{x})~|~\mathbf{x}\in\mathcal{C}\right\}=\frac{1}{2}\left(\mathbb{E}\left\{Y~|~\mathbf{x}\in\mathcal{T}_{k,1}(\mathcal{C})\right\}-\mathbb{E}\left\{Y~|~\mathbf{x}\in\mathcal{T}_{k,-1}(\mathcal{C})\right\}\right)=\Delta^{1/2}(k;\mathcal{C},\mathcal{D}_{n}).$ (H.19)